# ChartGalaxy: A Dataset for Infographic Chart Understanding and Generation

### v3

@@ Abstract
Infographic charts are a powerful medium for communicating abstract data by combining visual elements (e.g., charts, images) with textual information. However, their visual and structural richness poses challenges for large vision-language models (LVLMs), which are typically trained on plain charts. To bridge this gap, we introduce ChartGalaxy, a million-scale dataset designed to advance the understanding and generation of infographic charts. The dataset is constructed through an inductive process that identifies 75 chart types, 330 chart variations, and 68 layout templates from real infographic charts and uses them to create synthetic ones programmatically. We showcase the utility of this dataset through: 1) improving infographic chart understanding via fine-tuning, 2) benchmarking code generation for infographic charts, and 3) enabling example-based infographic chart generation. By capturing the visual and structural complexity of real design, ChartGalaxy provides a useful resource for enhancing multimodal reasoning and generation in LVLMs.
@@ 1 Introduction
Infographic charts are widely recognized as an effective form for communicating data and are commonly used in news media, business, and education [1, 2]. By integrating visual elements—charts, imagery such as pictograms and metaphorical graphics—with textual information, they present abstract data in a manner that is both engaging and easy to understand, making data more accessible to a broad audience. Despite their effectiveness for human audiences, foundation models, such as GPT-4 [3], Gemini [4], and LLaVA [5], face considerable challenges in automatically understanding infographic charts. The intricate interplay between visual and textual elements, diverse layout styles, and the need for cross-modal semantic reasoning pose significant difficulties. Moreover, automatically generating high-quality infographic charts also remains an open challenge. While human designers can create visually diverse and semantically rich infographic charts, this process is time-consuming and requires expertise. Meanwhile, AI-generated charts often suffer from issues such as low data fidelity, modest visual quality, limited diversity, and a lack of coherence across modalities. This highlights the critical need for a comprehensive dataset of infographic charts that can support the development of deep learning models capable of both automatic understanding and generation. However, existing efforts focus on constructing datasets that are mostly limited to plain charts, failing to capture the diverse range of design styles and layouts that are key characteristics of infographic charts. This limits the ability of the trained models to generalize across different real-world applications where infographic charts are commonly used.
@@
To address this limitation, we build ChartGalaxy, a million-scale dataset of high-quality real and synthetic infographic charts to facilitate automated understanding and generation. As shown in Fig. 2, we build ChartGalaxy in two steps: 1) collecting real infographic charts; 2) programmatically creating synthetic infographic charts. The real infographic charts are collected from 19 reputable chart-rich websites, such as Pinterest [6], Visual Capitalist [7], Statista [8], and Information is Beautiful [9]. The synthetic infographic charts are created following an inductive structuring process [10]. Specifically, we identify design patterns grounded in real infographic charts, including 75 chart types (e.g., bar charts), 330 chart variations that reflect different visual element styles, and 68 layout templates that define spatial relationships among elements. Based on these patterns, we then programmatically generate synthetic ones. The core of the generation is a human-in-the-loop pipeline that iteratively extracts and expands layout templates from real infographic charts using a detection model trained on synthetic infographic charts.
@@
The final ChartGalaxy dataset includes 1,151,087 programmatically created infographic charts and 104,519 real infographic charts. It is characterized by two key features. First, the high-quality infographic designs and associated templates from these reputable websites ensure a rich diversity in design styles and structural complexity. Second, each infographic chart, whether real or synthetic, is paired with the tabular data used to create it, enabling a clear mapping between the data and its visual representation. Together, these features make ChartGalaxy a useful dataset for training and evaluating LVLMs designed for the automatic understanding and generation of infographic charts. We demonstrate the utility of ChartGalaxy through three representative applications, each highlighting a distinct aspect of its value (Fig. 1). First, to evaluate and improve the ability of foundation models to understand infographic charts, we introduce a dataset for infographic chart understanding through the task of visual question answering (VQA). Second, to assess the capacity of models to generate executable representations of complex visual layouts, we present a benchmark for infographic chart code generation. Third, to explore the use of ChartGalaxy in creative content generation, we develop an example-based infographic chart generation method.
@@ 2 Related Work
Early efforts in chart dataset construction primarily focus on building collections of plain charts to support chart understanding and generation [11, 12]. These datasets can be further categorized into three types based on their sources: synthetic datasets, web-crawled datasets, and mixed datasets. Synthetic datasets are programmatically generated, using tabular data drawn from probability distributions [13, 14], crawled from online data sources [11, 15, 16, 17, 18, 19], or simulated from large language models [17, 20, 21, 22]. While this method enables large-scale dataset construction, the controlled generation process often results in a limited diversity of chart types and visual styles, which reduces generalizability to more varied real-world scenarios. To improve diversity, web-crawled datasets have been introduced, which collect charts from chart-sharing websites [23, 24, 25, 26, 27, 28, 29, 30], charting library galleries such as Matplotlib [31, 32], and publications in academic repositories such as ArXiv [33, 34, 35]. These datasets capture a broader variety of human-designed chart styles, but their overall size is often limited due to the time-consuming nature of manual verification and annotation. To overcome the scale limitations while preserving diversity, mixed datasets are then introduced. These datasets merge web-crawled charts with synthetically generated ones in a two-step workflow: 1) collect real charts from the internet and 2) manually synthesize additional charts that follow the same real-world design patterns [36, 37, 38, 39, 40]. This hybrid strategy ensures real-world coverage while increasing dataset size.
In contrast, infographic charts remain underrepresented in the aforementioned datasets. This creates challenges in evaluating and developing LVLMs’ capabilities on infographic charts [41]. To bridge this gap, recent efforts have focused on building specific datasets for infographic charts. InfographicVQA makes an initial effort by searching “infographics” on the internet and scraping 5,485 infographics [42]. More recently, ChartQAPro provides a more challenging benchmark comprising 190 infographic charts, 258 dashboards, and 893 plain charts [41]. However, these datasets are limited in scale. Moreover, synthesizing infographic charts is challenging, given the intricate interplay between visual and textual elements. To overcome these limitations, we develop an automatic infographic chart generation method that synthesizes infographic charts by leveraging the layout templates and chart variations extracted from real designs.
@@ 3.1 Method Overview
The chart collection stage aims to collect real infographic charts by crawling 19 chart-rich websites, such as Pinterest [6], Visual Capitalist [7], Statista [8]. The full list of these websites is provided in Supp. A. To ensure data quality, we remove duplicate images using perceptual hashing [43] and CLIP similarity [44]. We then extract the data table for each chart using GPT-4o-mini and Gemini-2.0-Flash independently, retaining only those with consistent tabular outputs. This results in 104,519 real infographic charts with corresponding tables. The chart creation stage follows an inductive structuring process that extracts design patterns, such as layout templates and chart variations, from real infographic charts and then uses these patterns to programmatically create high-quality synthetic charts. It includes three steps: 1) identifying chart types and their variations, 2) extracting layout templates, and 3) creating synthetic infographic charts with the identified design patterns.
@@ 3.2 Chart Type and Variation Identification
We first summarize 75 chart types observed in the collected real infographic charts, drawing on two existing taxonomies: Data Viz Project [45] and Datylon [46]. For each chart type, we extract chart variations featuring diverse visual styles, such as element shapes and icon placement. This results in 330 chart variations in total. The full lists of chart types and variations are provided in Supp. B.1 and B.2. We implement these chart types and variations using the expressive D3.js [47], which supports visual features unavailable in libraries like Matplotlib or Seaborn.
@@ 3.3 Layout Template Extraction
A layout template defines the spatial relationships among the text and visual elements in an infographic chart. Example templates are shown on the bottom-left corner of each chart in Fig.  3. We adopt a human-in-the-loop pipeline to initialize and expand these templates from real infographic charts.
Initialization   Three co-authors manually annotate the bounding boxes of the text, image, and chart in 1,500 real infographic charts sampled from two high-quality sources: Statista [8] (clean, minimalist designs) and Visual Capitalist [7] (denser and more pictorial designs). From these annotations, we summarize an initial set of 55 layout templates that capture elements’ relative positions (e.g., title on the top-left, chart on the bottom-right) and pairwise overlaps (overlapping or not).
Expansion   To ensure the coverage and diversity of templates, we build a detection model to analyze the unlabeled real infographic charts and systematically expand the template set. Using the initial templates, we programmatically create 120,000 synthetic infographic charts (Sec. 3.4), each with annotated bounding boxes. We then develop a detection model by fine-tuning InternImage [48] along with the DINO [49] detector on these synthetic charts. Applied to real unlabeled charts, this model detects chart and image regions [50], while text is extracted using PP-OCRv4 [51]. We then compare the detected layouts with the existing templates using LTSim [52], a state-of-the-art method for measuring layout similarity. Layouts with low similarity scores are flagged as potential new templates. Next, we cluster these layouts using k-means ($k=50$) and manually examine the cluster centroids to identify distinct layouts. This process yields 13 additional layout templates, expanding the set to 68 templates in total. The full list is provided in Supp. C.
@@ 3.4 Template-based Infographic Chart Creation
The creation process involves three steps: 1) curating data tables; 2) generating/recommending elements based on the data table; and 3) optimizing the layout based on the selected template.
Tabular data curation   To enhance data diversity for chart generation, we build a rich repository of real and synthetic tabular data. For real data, we collect 200,085 tables from well-established sources, including VizNet [53], UN Data [54], Our World in Data [55], and Papers with Code [56]. For synthetic data, we generate 98,483 tables with Gemini-2.0-Flash following Han et al.’s method [21]. To facilitate downstream processing, we also complement each table with a topic (e.g., “US election,” “NBA play-offs”) extracted by Gemini-2.0-Flash and several data facts (e.g., trends, comparisons) [57].
@@
Text. The titles and subtitles are generated using a retrieval-augmented prompting strategy to reflect real-world usage. Specifically, we use Sentence-BERT [58] to retrieve the three most relevant real infographic charts according to the data topics and data facts. Using these retrieved examples as in-context references, we prompt Gemini-2.0-Flash to generate titles and subtitles aligned with the tabular data.
@@
Image. To recommend images for infographic charts, we construct an image repository and retrieve semantically relevant images based on the associated tabular data. Images are selectively collected from publicly available resources such as Icon645 [59], Flaticon [60], and Noun Project [61]. We apply heuristic filters to remove low-quality images, excluding those with low ink ratios, poor resolution, or extreme aspect ratios. Using Gemini-2.0-Flash, we then generate descriptive keywords and captions for each image that depict its content and visual style. Images that are overly literal, visually cluttered, or lack symbolic clarity are discarded. This process results in a curated repository of 681,459 high-quality images. To retrieve relevant images for each infographic chart, we compute the similarity between the image keywords and the generated chart title using Sentence-BERT [58] embeddings.
@@
Chart. Chart generation proceeds in two steps: finding suitable chart types and rendering with a specific chart variation. Chart types are determined by analyzing the data attributes (e.g., categorical, numerical) and characteristics (e.g., scales). For example, a data table with one categorical column and two numerical columns may be mapped to a scatter plot. When multiple chart types are suitable, we prompt Gemini-2.0-Flash to select the optimal one based on data-chart compatibility [62]. A full list of mapping rules and prompts used for chart selection is provided in Supp. B.3. Once the chart type is selected, we choose a specific variation using an adaptive sampling strategy that favors underrepresented variations to ensure distributional balance. For variations requiring additional images (e.g., Figs. 3(a), 3(c) and 3(d)), we retrieve relevant ones from our curated image repository using the above image recommendation method. Finally, we apply semantically resonant color palettes for chart rendering based on the generated chart titles and subtitles. These palettes are extracted from the collected real infographic charts following Liu et al.’s method [63]. If a selected palette contains fewer colors than required, Chen et al.’s method is used to supplement additional harmonic and discriminable colors [64].
@@
Layout optimization   Previous research has shown that a compact layout with appropriate white space enhances both visual appeal and data clarity [65]. Consequently, we aim to select the template with the highest ink ratios while preserving readability. To this end, we first filter out templates that are incompatible with the generated elements (e.g., unintended overlaps between elements). For each remaining template, we initialize the positions for the set of elements, optimize the layout to reduce unnecessary white space, and compute the ink ratio. The template with the highest ink ratio is then selected for chart creation.
@@
Here, $e_{i}$ is the pixel set of an element, $f(e)$ denotes the pixel set within the tight-fitting bounding box of $e$, and $\mathcal{E}$ is the set of elements. We enforce that a minimum pairwise distance $d$ between element contours $\partial e_{i}$ and $\partial e_{j}$ is larger than a given threshold $p$. This layout optimization is formulated as a constrained packing problem and solved by grid search [66]. Fig. 3 shows representative examples of the synthetic charts. For more examples, please refer to Supp. C.

### v5

@@ Abstract
Infographic charts are a powerful medium for communicating abstract data by combining visual elements (e.g., charts, images) with textual information. However, their visual and structural richness poses challenges for large vision-language models (LVLMs), which are typically trained on plain charts. To bridge this gap, we introduce ChartGalaxy, a million-scale dataset designed to advance the understanding and generation of infographic charts. The dataset is constructed through an inductive process that identifies 75 chart types, 440 chart variations, and 68 layout templates from real infographic charts and uses them to create synthetic ones programmatically. We showcase the utility of this dataset through: 1) improving infographic chart understanding via fine-tuning, 2) benchmarking code generation for infographic charts, and 3) enabling example-based infographic chart generation. By capturing the visual and structural complexity of real design, ChartGalaxy provides a useful resource for enhancing multimodal reasoning and generation in LVLMs.
@@ 1 Introduction
Infographic charts are widely recognized as an effective form for communicating data and are commonly used in news media, business, and education (Cui et al., 2022; Elaldi & Çifçi, 2021). By integrating textual information and visual elements such as charts and images, they present abstract data in a manner that is both engaging and easy to understand, making data more accessible to a broad audience. Despite their effectiveness for human audiences, foundation models, such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and LLaVA (Liu et al., 2023c), face considerable challenges in automatically understanding infographic charts. The intricate interplay between visual and textual elements, diverse layout styles, and the need for cross-modal semantic reasoning pose significant difficulties. Moreover, automatically generating high-quality infographic charts remains an open challenge. While human designers can create visually diverse and semantically rich infographic charts, this process is time-consuming and requires expertise. Meanwhile, AI-generated charts often suffer from issues such as low data fidelity, modest visual quality, limited diversity, and a lack of coherence across modalities. This highlights the critical need for a comprehensive dataset of infographic charts that enables model development in both automatic understanding and generation. However, existing efforts focus on constructing datasets that are mostly limited to plain charts, failing to capture the diverse range of design styles and layouts that are key characteristics of infographic charts. This limits the ability of the trained models to generalize across different real-world applications where infographic charts are commonly used.
@@
To address this limitation, we build ChartGalaxy, a million-scale dataset of high-quality real and synthetic infographic charts to facilitate automated understanding and generation. As shown in Fig. 2, we build ChartGalaxy in two steps: 1) collecting real infographic charts; 2) programmatically creating synthetic infographic charts. The real infographic charts are collected from 18 reputable chart-rich websites, such as Visual Capitalist and Statista . The synthetic infographic charts are created following an inductive structuring process (Schadewitz & Jachna, 2007). Specifically, we identify design patterns grounded in real infographic charts, including 75 chart types (e.g., bar charts), 440 chart variations that reflect different visual element styles, and 68 layout templates that define spatial relationships among elements. Based on these patterns, we then programmatically generate synthetic ones. The core of the generation is a human-in-the-loop pipeline that iteratively extracts and expands layout templates from real infographic charts using a detection model trained on synthetic infographic charts.
@@
The final ChartGalaxy dataset includes 1,701,356 programmatically created infographic charts and 61,833 real infographic charts. It is characterized by two key features. First, the high-quality infographic designs and associated templates from these reputable websites ensure a rich diversity in design styles and structural complexity. Second, each infographic chart, whether real or synthetic, is paired with its source data table, enabling a clear mapping between data and its visual representation. Together, these make ChartGalaxy well-suited for training and evaluating LVLMs for automatic infographic understanding and generation. We demonstrate the utility of ChartGalaxy through three representative applications, each highlighting a distinct aspect of its value (Fig. 1). First, to evaluate and improve the ability of foundation models to understand infographic charts, we introduce a dataset for infographic chart understanding through the task of visual question answering (VQA). Second, to assess the capacity of models to generate executable representations of complex visual layouts, we present a benchmark for infographic chart code generation. Third, to explore the use of ChartGalaxy in creative content generation, we develop an example-based infographic chart generation method.
@@ 2 Related Work
Early efforts in chart dataset construction primarily focus on building collections of plain charts to support chart understanding and generation (Hu et al., 2024; Yang et al., 2024a). These datasets can be further categorized into three types based on their sources: synthetic datasets, web-collected datasets, and mixed datasets. Synthetic datasets are programmatically generated, using tabular data drawn from probability distributions (Kafle et al., 2018; Kahou et al., 2017), collected from online data sources (Hu et al., 2024; Methani et al., 2020; Chaudhry et al., 2020; Xu et al., 2023; Tang et al., 2023; Akhtar et al., 2023), or simulated from large language models (Xu et al., 2023; Xia et al., 2024; Han et al., 2023; Zhao et al., 2025). While this method enables large-scale dataset construction, the controlled generation process often results in a limited diversity of chart types and visual styles, which reduces generalizability to more varied real-world scenarios. To improve diversity, web-collected datasets have been introduced, which collect charts from chart-sharing websites (Savva et al., 2011; Choi et al., 2019; Masry et al., 2022; Kantharaj et al., 2022a; Huang et al., 2025; Obeid & Hoque, 2020; Kantharaj et al., 2022b; Masry et al., 2024), charting library galleries such as Matplotlib (Wu et al., 2024; Yang et al., 2024b), and publications in academic repositories such as ArXiv (Wang et al., 2024; Hsu et al., 2021; Zhang et al., 2024). These datasets capture a broader variety of human-designed chart styles, but their overall size is often limited due to the time-consuming nature of manual verification and annotation. To overcome the scale limitations while preserving diversity, mixed datasets are then introduced. These datasets merge web-collected charts with synthetically generated ones in a two-step workflow: 1) collect real charts from the internet and 2) manually synthesize additional charts that follow the same real-world design patterns (Liu et al., 2023a; b; Masry et al., 2023; Meng et al., 2024; Yang et al., 2025). This hybrid strategy ensures real-world coverage while increasing dataset size.
In contrast, infographic charts remain underrepresented in the aforementioned datasets. This creates challenges in evaluating and developing LVLMs’ capabilities on infographic charts (Masry et al., 2025). To bridge this gap, recent efforts have focused on building specific datasets for infographic charts. InfographicVQA makes an initial effort by searching “infographics” on the internet and scraping 5,485 infographics (Mathew et al., 2022). More recently, ChartQAPro provides a more challenging benchmark comprising 190 infographic charts, 258 dashboards, and 893 plain charts (Masry et al., 2025). However, these datasets are limited in scale. Moreover, synthesizing infographic charts is challenging, given the intricate interplay between visual and textual elements. To overcome these limitations, we develop an automatic infographic chart generation method that synthesizes infographic charts by leveraging the layout templates and chart variations extracted from real designs.
@@ 3.1 Method Overview
The real infographic charts are collected from two main sources. First, we gather infographics from 18 well-known chart-rich websites that permit research use, such as Visual Capitalist and Statista . The full list of these websites and their licenses is provided in Appendix A. Second, to further increase the diversity of infographic styles, we retrieve infographic charts via Google Images and Bing Images, following prior work (Masry et al., 2024). We apply the platforms’ built-in license filters to retain only images available under Creative Commons licenses. The ethical consideration in the data collection process is further discussed in Ethics Statement. To ensure data quality, we remove duplicate images using Perceptual Hashing and CLIP similarity (Radford et al., 2021). Moreover, we extract per-chart tabular data using LVLMs in a multi-step, human-in-the-loop verification pipeline. This pipeline ensures accuracy and reduces model-induced noise. Details of this verification pipeline are discussed in Appendix B. This results in 61,833 real infographic charts with corresponding tables. The synthetic infographic chart creation stage follows an inductive structuring process that extracts design patterns, such as layout templates and chart variations, from real infographic charts and then uses these patterns to programmatically create high-quality synthetic charts. It includes three main steps: 1) identifying chart types and their variations, 2) extracting layout templates, and 3) creating synthetic infographic charts.
@@ 3.2 Chart Type and Variation Identification
We first summarize 75 chart types observed in the collected real infographic charts, drawing on two existing taxonomies: Data Viz Project and Datylon . For each chart type, we extract chart variations featuring diverse visual styles, such as element shapes and icon placement. This results in 440 chart variations in total. The full lists of chart types and variations are provided in Appendix D.1 and D.2. We implement these chart types and variations using the expressive D3.js (Bostock et al., 2011), which supports visual features unavailable in libraries like Matplotlib or Seaborn.
@@ 3.3 Layout Template Extraction
A layout template defines the spatial relationships among the text and visual elements in infographic charts. Example templates are shown in the bottom-left corner of each chart in Fig. 3. We adopt a human-in-the-loop pipeline to initialize and expand these templates from real infographic charts.
Initialization Three co-authors manually annotate the bounding boxes of the text, images, and charts in 1,500 real infographic charts sampled from two high-quality sources: Statista (clean, minimalist designs) and Visual Capitalist (denser and more pictorial designs). From these annotations, we summarize an initial set of 55 layout templates that capture elements’ relative positions (e.g., title on the top-left, chart on the bottom-right) and pairwise overlaps (overlapping or not).
Expansion To ensure the coverage and diversity of templates, we build a detection model to analyze the unlabeled real infographic charts and expand the layout template set. Using the initial templates, we programmatically create 120,000 synthetic infographic charts (Sec. 3.4), each with annotated bounding boxes. We then develop a detection model by fine-tuning InternImage (Wang et al., 2023) along with the DINO (Zhang et al., 2023) detector on these synthetic charts. We use this model to detect chart and image regions (Zhu et al., 2025a) and use PP-OCRv4 to extract text. We then compare the detected layouts with the existing templates using LTSim (Otani et al., 2024), a state-of-the-art method for measuring layout similarity. Layouts with low similarity scores are flagged as potential new templates. Next, we cluster these layouts using k-means ($k=50$) and manually examine the cluster centroids to identify distinct layouts. This process yields 13 additional layout templates, expanding the set to 68 templates in total. The full list is provided in Appendix F.
@@ 3.4 Template-based Infographic Chart Creation
The creation process involves three steps: 1) curating data tables; 2) generating/recommending elements based on the data table; and 3) optimizing the layout based on the selected template. Fig. 3 shows six examples of the synthetic charts. More examples are provided in Appendix G.
Tabular data curation To enhance data diversity for chart generation, we build a rich repository of real and synthetic tabular data. For real data, we collect 200,085 tables from well-established sources, including VizNet (Hu et al., 2019)UN data Our World in Data , and Papers with Code . For synthetic data, we generate 98,483 tables with Gemini-2.0-Flash following Han et al. (2023). To facilitate downstream processing, we also complement each table with a topic (e.g., “NBA play-offs”) extracted by Gemini-2.0-Flash and several data facts (e.g., trends, comparisons) following Wang et al. (2020).
@@
Text. The title and subtitle are generated using a retrieval-augmented prompting strategy. Specifically, we use Sentence-BERT (Reimers & Gurevych, 2019) to retrieve the three most relevant real infographic charts according to the data topic and data fact. Using them as in-context references, we prompt Gemini-2.0-Flash to generate the title and subtitle aligned with the data table.
@@
Image. To recommend images for infographic charts, we build a repository and retrieve semantically relevant images from it using the associated data tables. Images are collected from publicly available icon resources such as Icon645 (Lu et al., 2021) and Noun Project . We apply heuristic filters to remove low-quality images, excluding those with low ink ratios, poor resolution, or extreme aspect ratios. Using Gemini-2.0-Flash, we then generate descriptive keywords and captions for each image that depict its content and visual style. Images that are overly literal, visually cluttered, or lack symbolic clarity are discarded. This process results in a curated repository of 681,459 images. To retrieve relevant images for each infographic chart, we compute the similarity between the image keywords and the generated chart title using Sentence-BERT embeddings (Reimers & Gurevych, 2019).
@@
Chart. Chart generation proceeds in two steps: selecting a suitable chart type and rendering with a specific chart variation. First, based on the attributes (e.g., categorical, numerical) and characteristics (e.g., scales) of the given data table, we identify candidate chart types following the predefined data-to-chart mapping rules. For example, a data table with one categorical column and two numerical columns may be mapped to a scatter plot. When multiple chart types are suitable, we prompt Gemini-2.0-Flash to select the optimal one, considering data types, value distributions, and temporal patterns (Luo et al., 2018). A full list of mapping rules and prompts is provided in Appendix E. Next, we choose a specific variation under the selected chart type using adaptive sampling (Goodfellow et al., 2016) that favors underrepresented variations to maintain distributional balance. For variations requiring additional images (e.g., Fig. 3(a)), we retrieve the relevant ones from our curated image repository. Finally, we apply semantically resonant color palettes for chart rendering based on the generated chart titles and subtitles. These palettes are extracted from the collected real infographic charts (Liu et al., 2024). If a selected palette contains fewer colors than required, we supplement additional harmonic and discriminable colors (Chen et al., 2025).
@@
Layout optimization Previous research has shown that a compact layout with appropriate white space enhances both visual appeal and data clarity (Coursaris & Kripintris, 2012). Consequently, we aim to select the template with the highest ink ratios while preserving readability. To this end, we first filter out templates that are incompatible with the generated elements (e.g., unintended overlaps between elements). For each remaining template, we initialize the positions for the set of elements, optimize the layout to reduce unnecessary white space, and compute the ink ratio. The template with the highest ink ratio is then selected for chart creation.
@@
Here, $e_{i}$ is the pixel set of an element, $f(e)$ denotes the pixel set within the tight-fitting bounding box of $e$, and $\mathcal{E}$ is the set of elements. We enforce that a minimum pairwise distance $d$ between element contours $\partial e_{i}$ and $\partial e_{j}$ is larger than a given threshold $p$. This layout optimization is formulated as a constrained packing problem and solved by grid search (Lodi et al., 1999).